write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b
(-3,6);4x=9y+8
The equation of the line is Y=

Answer by mananth(16949)

4x=9y+8
9y=4x-8
y= 4/9 x -8/9
slope of line = 4/9
The parallel line will also have slope of 4/9. It passes through -3 6
Use the equation y-y1 / x-x1= m
y-6 = 4/9(x-(-3))
y-6= 4/9 (x+3)
9y-54= 4x+12
9y= 4x+12+54
9y= 4x +66
y= 4/9 x +66/9